5.) Determine the present value if $15,000 is to be received at the end of eight years and the discount rate is 9 percent. How would your answer
change if you had to wait six years to receive the$15,000?


13.) a.) what would be the future value of $7,455 invested annually for nine years beginning one year from now if the annual interest rate in 9 percent





6.) assume you are planning to invest $5,000 each year for six years and will earn 10 percent per year. Determine the future valuye of this annuity if your first $5,000 is invested at end of this year.


7.)what is the present value of a loan that calls for the payment of $500 per year for six years if the discount rate is 10 percent and the first payment will be made one year from no? how would your answer change if the $500 per year occurred for ten years?

8.) determine the annual payment on a $500,000, 12 percent business loan from a commercial bank that is to be amortized over a 5 year period.





Answer by rfer(16322)

5)
P=A/(1+r/n)^nt
P=15000/(1.09)^8
P=15000/(1.992562642)
P=$7528
-----------------------
P=15000/(1.09)^6
P=15000/1.6771
P=$8940
-----------------------
13)
FV=P(1+r)^9
FV=7455(1.09)^9
FV=7455*2.171893279
FV=$16191
-----------------------
6)
FV=(P((1+i)^n-1)/i
FV=(5000((1.1)^6-1)/.1
FV=(5000((1.771561)-1)/.1
FV=3857.81/.1
FV=$38578.05
---------------